# Nef Cones of the Hilbert Schemes of Points on Generalized Cayley K3 Surfaces

### v1

@@ 1. Introduction
Bridgeland stability conditions provide a new conceptual approach to studying these problems. Bayer and Macrì first used wall-crossing to describe the nef and movable cones of moduli spaces of sheaves on K3 surfaces [5], and subsequently established the Positivity Lemma [4], which produces nef divisors from stability conditions and identifies the curves orthogonal to them. Building on these results, the wall–chamber structure for Hilbert schemes of points was analyzed in detail [8].
@@ Acknowledgments
The author would like to thank Yongnam Lee for detailed comments on the draft of this paper. This work was supported by the Institute for Basic Science (IBS-R032-D1).
@@ 2.5. Bridgeland stability conditions
These divisors are closely connected to the birational geometry of the moduli space. If a curve has intersection number zero with $D_{\sigma}$, then it is contracted by the morphism associated to $D_{\sigma}$. Such contractions arise when the stability condition lies on a wall, linking wall-crossing phenomena with the structure of the nef and Mori cones of $S^{[n]}$.
In the case of the Hilbert scheme $X=S^{[n]}$, we take the Mukai vector v=(1,0,1n)v=(1,0,1-n), for which $M_{\sigma}(v)\cong S^{[n]}$. In this setting, one can describe the walls in the $(H,D)$-slice explicitly; the following results from [8] give formulas for the numerical wall and the Gieseker wall.
@@
Proposition 2.9 ([8, Prop. 3.8]).

### v2

@@ 1. Introduction
Bridgeland stability conditions provide a new conceptual approach to studying these problems. Bayer and Macrì first used wall-crossing to describe the nef and movable cones of moduli spaces of sheaves on K3 surfaces [5], and subsequently established the Positivity Lemma [4], which produces nef divisors from stability conditions and identifies the curves orthogonal to them. Liu [20] gave a general formulation of the local Bayer–Macrì decomposition of nef divisors. The wall–chamber structure for Hilbert schemes of points was analyzed in detail in [8].
@@ Acknowledgments
The author would like to thank Yongnam Lee for detailed comments on the draft of this paper, and Wanmin Liu for helpful comments on the references. This work was supported by the Institute for Basic Science (IBS-R032-D1).
@@ 2.5. Bridgeland stability conditions
These divisors are closely connected to the birational geometry of the moduli space. If a curve has intersection number zero with $D_{\sigma}$, then it is contracted by the morphism associated to $D_{\sigma}$. Such contractions arise when the stability condition lies on a wall, linking wall-crossing phenomena with the structure of the nef and Mori cones of $M_{\sigma}(v)$.
We begin by recalling that a general description of Bayer–Macrì line bundles on Bridgeland moduli spaces over surfaces is provided by the local Bayer–Macrì decomposition established by Liu [20]. In the present paper, we focus on the case of the Hilbert scheme $X=S^{[n]}$, corresponding to the Mukai vector v=(1,0,1n)v=(1,0,1-n); for later use, we adopt the description specialized to this setting, which appeared later in [8].
@@
Proposition 2.9 (cf. [20, Lemma 4.8]; see also [8, Prop. 3.8]).